I need help in algebra in functions and notation.Solve G(x) for the given domain. G(x)=3x exponent 2-2x-1.
G(a+b)=


Please help me with this algebra and explain to me how to find the answers.Thank you for your time.

Answer by stanbon(75887)

John tosses a coin three times in succession. He wins $1 for head on the first toss, $2 for head on the second toss and $3 for head on the third toss. For each tail he losses $1.5. Let X be the random variable of John's winning. Find distinct values of X and their probability distributions
I started to do this problem but I am not sure that I am using the correct formula. Could you please show me how to get the solution.
----------------------
Results are
hhh----$6-----1/8--------------6*1/8 = 12/16
hht----$1.5 1/8----------1.5*(1/8) = 3/16
hth----$2.5---1/8----------2.5*(1/8) = 5/16
thh----$3.5---1/8----------3.5*(1/8) = 7/16
tth----$ 0 ---1/8------------0*(1/8) = 0/16
tht----$ -1---1/8---------- -1*(1/8) = -1/8
htt----$ -2---1/8---------- -2*(1/8) = -4/16
ttt----$-4.5--1/8-------- -4.5*(1/8) = -9/16
==================================================
Expected Value = $13/16 = 81.25 cents
==================================================
Cheers,
Stan H.